### v1

# Can You Trust LLM Judgments? Reliability of LLM-as-a-Judge

Abstract
Large Language Models (LLMs) have become increasingly powerful and ubiquitous, but their stochastic nature poses challenges to the reliability of their outputs. While deterministic settings can improve consistency, they do not guarantee reliability, as a single sample from the model’s probability distribution can still be misleading. Building upon the concept of LLM-as-a-judge, we introduce a novel framework for rigorously evaluating the reliability of LLM judgments, leveraging McDonald’s omega. We evaluate the reliability of LLMs when judging the outputs of other LLMs on standard single-turn and multi-turn benchmarks, simultaneously investigating the impact of temperature on reliability. By analyzing these results, we demonstrate the limitations of fixed randomness and the importance of considering multiple samples, which we show has significant implications for downstream applications. Our findings highlight the need for a nuanced understanding of LLM reliability and the potential risks associated with over-reliance on single-shot evaluations. This work provides a crucial step towards building more trustworthy and reliable LLM-based systems and applications.
Can You Trust LLM Judgments? Reliability of LLM-as-a-Judge
Kayla Schroeder Department of Statistics Northwestern University kaylaschroeder2026@u.northwestern.edu                        Zach Wood-Doughty Department of Computer Science Northwestern University zach@northwestern.edu
1 Introduction
In recent years, Large Language Models (LLMs) have experienced rapid advancements and widespread adoption. Such growth has been embraced widely, with applications ranging from marketing to biotechnology to music (Gozalo-Brizuela and Garrido-Merchán, 2023). As reliance upon LLMs grows, ensuring the quality and reliability of their outputs becomes crucial.
A singular output from an LLM is the result of a draw from a probabilistic distribution (Vaswani, 2017). Parameters that vary the resulting draw, namely temperature and top-k, are increased to obtain more novel or "creative" results - a desirable quality of LLMs (Peeperkorn et al., 2024). However, this inherent stochastic nature raises concerns about the reliability of individual outputs. Given that an output from an LLM is only a single instantiation of the model, reliability of LLM output is of utmost importance (Mittelstadt et al., 2023). Consider a scenario where an LLM is tasked with evaluating the quality of another LLM that’s handling social media moderation. In this high-stakes context, ensuring the reliability of LLM judgments is crucial to user safety. If an LLM’s judgment is influenced by random chance rather than its underlying capabilities, it could lead to serious consequences, such as the proliferation of harmful content.
Many research works aim to circumvent this issue by setting a fixed instantiation and using deterministic settings for the temperature and top-k parameters (Ouyang et al., 2023; Wei et al., 2024; Atil et al., 2024). These studies argue that if an LLM consistently produces the same output under these conditions, it can be considered reliable. However, consistent replication does not guarantee the reliability of the generated text.
Even with deterministic settings, a single LLM output remains a sample from the model’s probability distribution, subject to inherent randomness. This results in "fixed randomness," which can lead to significant limitations (Hellrich and Hahn, 2016). Consider our prior example of an LLM tasked with evaluating the quality of another LLM handling social media moderation. If the judging LLM fails to detect poor performance in a single instantiation due to randomness, grave repercussions can result. By assessing the reliability of LLM judgments, we can determine when additional human oversight is necessary to ensure the safety and accuracy of LLM-powered systems, especially as these systems evolve over time.
The ability to trust LLM outputs is of further importance as LLMs are increasingly used as evaluators (Desmond et al., 2024; Chan et al., 2023a). The LLM-as-a-judge paradigm, coined by Zheng et al. (2023), relies upon the internal consistency reliability of LLM judges. Widespread utilization of LLMs as judges has garnered concern. The ACL Rolling Review Reviewer Guidelines highlight this issue, voicing explicit concern about LLM-only evaluation without validation by listing the methodological issue "If LLMs are used as automated evaluators, is their reliability in this context sufficiently validated?" [^1] As of December 2024 at https://aclrollingreview.org/reviewerguidelines. This underscores the critical need for trust in LLM outputs and their reliability. Reliability here, however, is not explicitly defined, and inconsistencies in how reliability of LLMs should be quantified are present throughout the literature.
While some research has focused on determining LLM reliability by comparing outputs to ground truth labels (Zhou et al., 2024; Fu et al., 2023a), this approach is often limited by the availability of accurate and comprehensive ground truth data. As LLMs are more readily used in more complex and subjective contexts, such as in the scenario of an LLM tasked with evaluating the quality of another LLM handling social media moderation, existence of a ground truth response becomes increasingly impractical. Other studies have explored inter-rater reliability (Kollitsch et al., 2024; Wang et al., 2024), but this metric is not suitable for evaluating LLM judges, as it assumes multiple human evaluators rather than a single, potentially biased, LLM model. Additionally, inter-rater reliability does not take into account the additional nuance of the LLM judge being a model itself.
The unique characteristics of LLMs as evaluators necessitate a more nuanced approach to assessing their reliability. We need to consider factors such as the consistency of their judgments, their sensitivity to different prompts and datasets, and their ability to generalize to new domains. A more robust framework for assessing LLM-based evaluation is essential to ensure quality and fairness in AI-driven decision-making. A key component of this framework is a measure of the internal consistency reliability of LLM judgments.
Internal consistency reliability, a well-established statistical concept with a rich literature, can be effectively measured using metrics like Cronbach’s alpha and McDonald’s omega (Cronbach, 1951; McDonald, 2013; Malkewitz et al., 2023). These metrics can effectively quantify reliability of a model’s judgments by assessing the consistency of its responses to different prompts or tasks.
We introduce an innovative framework for evaluating the reliability of LLMs as judges. We illustrate that McDonald’s omega can effectively quantify the internal consistency of LLM judgments across replications (different runs where only the instantiation changes). Applying this to both single-turn and multi-turn tasks across varying temperatures, we investigate the impact of LLM-as-a-judge reliability. Further, we demonstrate the implications of reliability on downstream applications.  [^2] The code for the framework and analyses can be accessed at https://github.com/kaylaschroeder/llm_reliability.
2 Related Work
Reliability is a cornerstone of NLP research, particularly when human judgment is involved. Studies utilizing human labelers have long emphasized inter-annotator agreement (Bhowmick et al., 2008; Nowak and Rüger, 2010; Amidei et al., 2019). Similarly, topic modeling has extensively explored reliability through various methods, including similarity measures, statistical techniques, and domain expertise (Rieger et al., 2024; Schroeder and Wood-Doughty, 2024; Chuang et al., 2015).
Internal consistency reliability itself, especially Cronbach’s alpha, has been heavily utilized as an assessor of question quality in LLM contexts (Shang et al., 2024; Chan et al., 2023b; Biri et al., 2023). Bhandari et al. (2024) uses Cronbach’s alpha to compare LLM-generated questions with human-authored ones; Shi et al. (2023) uses Cronbach’s alpha to assess the reliability of questions designed to evaluate LLMs.
The concept of reliability appears widely in the LLM literature, but is not consistently defined. Some studies equate reliability with human-model agreement, requiring costly human evaluations (Sun et al., 2024; Zhang et al., 2024). Others rely on accuracy metrics, which necessitate ground truth labels (Zheng et al., 2024; Macherla et al., 2023; Wei et al., 2024), a requirement that proves increasingly unrealistic as the judgment task increases in complexity and subjectivity. Semantic consistency has also been proposed as a reliability measure (Govil et al., 2024; Raj et al., 2023), despite its lack of theoretical foundation and potential pitfalls, as detailed in Schroeder and Wood-Doughty (2024). Many studies also use the term "reliability" without rigorous definition (Tan et al., 2024; Liao et al., 2023; Gupta et al., 2024).
Currently, the LLM-as-a-judge literature predominantly relies upon accuracy and bias metrics to assess reliability (Chen et al., 2024; Wei et al., 2024; Ye et al., 2024). Some researchers advocate for combining responses from different LLM judgment models to improve reliability (Gu et al., 2024; Patel et al., 2024). To quantify reliability, human judgments are often used as a benchmark, with methods relying upon agreement with human evaluation and correlation with human-annotated error scores (Jung et al., 2024; Fu et al., 2023b; Dong et al., 2024). Our work pioneers the investigation of internal consistency reliability in the LLM-as-a-judge paradigm, addressing a critical gap in the literature.
3 Internal Consistency Reliability
Internal consistency reliability is a cornerstone of psychometric analysis and has been extensively studied. Cronbach’s alpha, introduced by Cronbach (1951), has been widely used to assess internal consistency, but relies on several assumptions that may not always hold (Malkewitz et al., 2023; Agbo, 2010). As proposed by McDonald (2013), McDonald’s omega addresses these limitations by accounting for a broader range of factor structures. Ultimately, omega provides a more robust measure of internal consistency and has been shown to be more robust to deviations from assumptions than alpha (Stensen and Lydersen, 2022). In this work, we adopt McDonald’s omega as our primary metric, recognizing its superior quantification of reliability.
3.1 McDonald’s Omega
McDonald’s omega ($\omega$) is formulated as
$\dfrac{\left(\sum_{i=1}^{n}\lambda_{i}\right)^{2}}{\left(\sum_{i=1}^{n}\lambda_{i}\right)^{2}+\sum_{i=1}^{n}\theta_{ii}+2\sum_{1\leq i<j\leq n}\theta_{ij}}$
where (in this setting) $n$ is the number of replications, $\lambda_{i}$ is the factor loading of the $i$th replication, $\theta_{ii}$ is the standardized error variance of the $i$th replication, and $\theta_{ij}$ is the covariance between replications $i$ and $j$ (McDonald, 2013). A replication in this context refers to an LLM judgment output that differs only in its instantiation from other replications for the same judgment prompt. Omega relies upon the common multidimensional methodology of factor analysis, of which the factor loadings are derived (Kline, 2014). Factor loadings are statistical measures derived from factor analysis. They quantify the strength of the relationship between each replication’s result and the underlying latent trait, which is the unobserved true judgment that the replications are intended to measure. In simpler terms, factor loadings indicate how much each replication contributes to our understanding of the overall judgment being measured. A higher factor loading suggests that the replication is a strong indicator of the latent trait.
To apply omega, we assume:
  1. 1.
    Additive measurement error: Each replication’s values are subject to random errors that are additive to the true score.
  2. 2.
    Uncorrelated errors: The errors associated with different replications are independent of each other.
  3. 3.
    Single latent trait: Each replication measures a single underlying trait.
Assumption 1, requiring additive measurement error, is a common assumption in statistical test theory, especially for items with similar response formats, making it a reasonable assumption in our context (Alagumalai and Curtis, 2005; Wadkar et al., 2016; Liu et al., 2010). Additionally, assumptions 2 and 3 are trivial in our setting as each replication is independent so errors are guaranteed to be uncorrelated, and each replication result is an outputted response to a singular prompt so it logically follows that each replication measures a single underlying construct.
4 Experiments and Data
We evaluate the reliability of the LLM-as-a-judge paradigm across diverse question formats and difficulty levels. Using LLM responses to benchmark questions, judgment LLMs are prompted repeatedly to select the "best" response based on factors like accuracy, utility, and relevance. We ran each judgment prompt 100 times, varying only the instantiation while holding other factors constant. Additionally, we varied the temperature parameter to further explore the influence of fixed randomness on LLM outputs. All judgment results were then assessed to uncover reliability by applying McDonald’s omega.
To obtain data for judgments, we utilize the widely accepted benchmarks BIG-Bench Hard (BBH) from Suzgun et al. (2022), SQuAD from Rajpurkar (2016), and MT-Bench from Zheng et al. (2023). These benchmarks offer two key advantages. First, they present diverse question formats: BBH and SQuAD questions are single-turn with specific target answers, while MT-Bench questions are multi-turn and open-ended. Second, they provide varying levels of difficulty, as evidenced by the significantly different accuracy rates of LLMs on BBH and SQuAD questions shown in Table 3. This diversity allows us to assess judgments under a range of conditions.
We first randomly selected one question from each category within each benchmark, for a total of 55 questions. Next, we generated five responses for each question using Chain-of-Thought prompting, as illustrated in Figure 1. The variation in prompting by benchmark is a result of the structuring of each benchmark; BBH responses require a single question, SQuAD responses require a single question as well as context surrounding the question, and MT-Bench requires a two-turn dialogue. Responses to questions are generated from five highly popular open-source LLMs: LLaMA-3-8B, Vicuna-7B-v1.5, Gemma-7B, Phi-2, and Falcon-7b (Dubey et al., 2024; Chiang et al., 2023; Team et al., 2024a; Javaheripi et al., 2023; Almazrouei et al., 2023). These models were chosen due to their popularity, with over 100,000 downloads in the prior month on Hugging Face. [^3] As of November 2024 at https://huggingface.co/models The generated responses are obtained using top-k of 50 and a temperature of 0.75.
Include step-by-step reasoning in answering the following question: [Question]
(a) Prompt template utilized to generate responses to BBH questions.
Include step-by-step reasoning in answering the following question: Context: [Context] Question: [Question] Answer:
(b) Prompt template utilized to generate responses to SQuAD questions.
Include step-by-step reasoning in answering the following question: [Question 1] [Response 1] [Question 2]
(c) Prompt template utilized to generate responses to the second turn and final of MT-Bench questions. To generate responses from the first turn, the template is identical to the BBH prompt.
Figure 1: Prompt templates utilized to generate responses from each of the varying benchmarks.
The model responses are then combined in a single prompt, as shown in Figure 2. To prevent judgment bias, the five responses to the respective benchmark question are included in the prompt in a random order as response options [A] - [E]. This random order was held constant across judgment replications, but shuffled for each new benchmark question judgment prompt. In addition to the five responses, the judgment prompt includes the original benchmark question (and context or previous turns if relevant) and prompts the LLM for a judgment of best response including Chain-of-Thought. Here, we utilize Chain-of-Thought from Wei et al. (2022) as it has been shown to dramatically improve LLM performance Feng et al. (2024); Suzgun et al. (2022); Chu et al. (2023). For the multi-turn responses from MT-Bench, options [A]-[E] were explicitly labeled as responses to the final turn, as shown in Figure 2(b).
You are a fair and objective judge tasked with selecting the strongest of the following responses to the provided question. Base your judgment upon the accuracy, utility, and relevance of each. Do not consider length of response, positioning of response or title of response in your judgment. Output the letter of the best response followed by an explanation, and strictly follow the following format: "Best Response: [[letter]]". Question: [Question] Responses: [A]: [Response A] [B]: [Response B] [C]: [Response C] [D]: [Response D] [E]: [Response E]
(a) Prompt template utilized to generate judgments of responses to BBH and SQuAD questions.
You are a fair and objective judge tasked with selecting the strongest of the following responses to the second provided question ("Question 2"). Question 1 is only provided for context. Base your judgment upon the accuracy, utility, and relevance of each. Do not consider length of response, positioning of response or title of response in your judgment. Output the letter of the best response followed by an explanation, and strictly follow the following format: "Best Response: [[letter]]". Question 1: [Question 1] Question 2: [Question 2] Responses: [A]: [Response A] [B]: [Response B] [C]: [Response C] [D]: [Response D] [E]: [Response E]
(b) Prompt template utilized to generate judgments of responses to MT-Bench questions.
Figure 2: Prompt templates utilized to generate judgments of responses to each of the varying benchmarks.
Each LLM utilized for judgments was prompted 100 times, varying only the instantiation, to judge which of the responses to the given benchmark question is best out of the five responses provided. Given the employment of Chain-of-Thought in the initial responses to the question, a single best response should exist in response options [A] - [E] for a judge to uncover because each of the five responses contains reasoning in addition to the question response. Similar logic is employed in Zheng et al. (2023) in assessment of judges as well. The process is then repeated, varying the temperature of the judgments with temperatures of 1, 0.75, 0.5, 0.25, and 0 (machine epsilon) while holding top-k constant at 50. Temperature has long been known to impact the "novelty" of LLM outputs (Peeperkorn et al., 2024; Davis et al., 2024). Varying the temperature provides a window into the variability across instantiations and allows for assessment of the impact of fixed randomness.
Selecting models that are high performers in the Chatbot Arena (Chiang et al., 2024), Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct are employed for the judgment task Zhu et al. (2023a); Team et al. (2024b); AI@Meta (2024). These selected models are all below the threshold of 10B parameters and are top performing models, ranked in the top 100, with rankings of 65 (Meta-Llama-3-8B-Instruct), 78 (Starling-LM-7B-beta), and 98 (Gemma-1.1-7b-it). In addition, we note that the selected models differ from those used for question answering in the prior step. This distinction is crucial to avoid circular reasoning, where LLMs would be judging responses they themselves generated. This approach also helps mitigate potential bias.
Reliability of these judgments are then calculated using McDonald’s omega and employing the reliabiliPy package (Fernández, 2022). It is important to note that in calculating the reliability, judgments that did not provide a judgment of best response were treated as belonging to the same category, thereby resulting in the calculated reliability serving as an upper bound on reliability. This is further explored in the experiment results. Results are obtained using a pair of Quadro RTX 8000 GPUs with 48GB of memory each and the transformers library (Wolf et al., 2020).
5 Experimental Results
Temperature BBH SQuAD MT-Bench
1 0.702 0.632 0.462
0.75 0.713 0.639 0.618
0.5 0.703 0.644 0.476
0.25 0.677 0.598 0.64
0 1 1 1
(a) Omega reliability of Starling-LM-7B-beta judgments.
Temperature BBH SQuAD MT-Bench
1 0.712 0.632 0.59
0.75 0.698 0.655 0.556
0.5 0.680 0.554 0.602
0.25 0.661 0.533 0.421
0 1 1 1
(b) Omega reliability of Meta-Llama-3-8B-Instruct judgments.
Temperature BBH SQuAD MT-Bench
1 0.723 0.64 0.605
0.75 0.77 0.73 0.585
0.5 0.788 0.751 0.732
0.25 0.803 0.77 0.637
0 1 1 1
(c) Omega reliability of Gemma-1.1-7b-it judgments.
Table 1: Omega reliability by LLM judge across temperatures and benchmarks.
Table 1 presents a comprehensive overview of judgment reliability for temperature variations across various benchmarks and LLMs. While one might expect that decreasing temperature would lead to a corresponding increase in reliability, as observed with a temperature of 0 (resulting in identical outputs), the reality is more intricate.
Our findings reveal a complex interplay between temperature and reliability. While some cases, such as Gemma-1.1-7b-it judgments of SQuAD responses (Table 1(c)), exhibit a clear decrease in reliability with increasing temperature, others, like Meta-Llama-3-8B-Instruct judgments of BBH prompts (Table 1(b)), show the opposite trend. This unexpected behavior highlights the nuanced relationship between temperature and LLM output quality. While higher temperatures can indeed lead to more novel and diverse outputs, they can also introduce variability and potential degradation in reliability. The ability to maintain reliability while increasing temperature is a promising avenue for future research, as it opens up opportunities to explore methods that generate novel, reliable, and high-quality outputs.
Reliability Interpretation
$\mathbf{\boldsymbol{\alpha}\boldsymbol{>}0.9}$ Excellent
$\mathbf{0.9>\boldsymbol{\alpha}>0.8}$ Good
$\mathbf{0.8>\boldsymbol{\alpha}>0.7}$ Acceptable
$\mathbf{0.7>\boldsymbol{\alpha}>0.6}$ Questionable
$\mathbf{0.6>\boldsymbol{\alpha}>0.5}$ Poor
$\boldsymbol{\alpha}\mathbf{<0.5}$ Unacceptable
Table 2: Rule of thumb for interpreting reliability measures.
While the reliability across temperatures shows promise, the reliability values themselves in Table 1 are a cause of concern, highlighting significant inconsistencies in judgments across replications. It’s important to note that the calculated reliability values are an optimistic estimate of the true reliability, as non-response judgments were treated as belonging to the same category. This assumption likely overestimates the reliability, particularly for models with high rates of non-response. Adhering to the widely accepted rule of thumb in Table 2 for interpreting reliability values in practice, we assess the reliability of our judgments (Gliem et al., 2003). Starling-LM-7B-beta and Meta-Llama-3-8B-Instruct exhibit questionable, at best, reliability across all benchmarks, with SQuAD and MT-Bench results consistently falling below the acceptable range (Tables 1(a) and 1(b), respectively). This low reliability raises concerns about the generalizability of their judgments. Gemma-1.1-7b-it fares better, with BBH judgments falling within the acceptable range or higher across all temperatures. SQuAD results are predominantly within the acceptable range, and no MT-Bench results fall within the unacceptable range (Table 1(c)). This makes Gemma-1.1-7b-it our most reliable judge, followed by Starling-LM-7B-beta and then Meta-Llama-3-8B-Instruct. Interestingly, this reliability ranking is the inverse of the models’ performance on the Chatbot Arena. This suggests a potential trade-off between model performance and judgment reliability, where models prioritizing performance may sacrifice reliability in their assessments.
Min Q1 Q2 Q3 Max
BBH 0 0 0 1 2
SQuAD 3 5 5 5 5
Table 3: Five-number summary of accuracies of responses [A]-[E] across all sampled questions for each benchmark.
To accurately assess benchmark performance, we must consider the inherent quality of response options [A]-[E]. As shown in Table 3, SQuAD questions exhibit significantly and consistently higher response accuracy compared to BBH questions. This distinction is pivotal. Lower response accuracy makes it easier to identify a clear best answer, as the difference in quality between options is more pronounced. In contrast, higher accuracy often results in multiple correct options, demonstrating judges’ difficulty in discerning the best response based solely on explanations. Consequently, the lower reliability observed for SQuAD across models and temperatures in Table 1 likely reflects this inherent difficulty in judging among highly accurate responses. This result is important to take into consideration in the employment of LLM-as-a-judge, as it points to potential struggle of LLM judges to differentiate between subtly nuanced correct responses.
Table 1 also shows that reliability is consistently worse for MT-Bench than for BBH and often worse for MT-Bench compared to SQuAD. Accuracies for MT-Bench are not provided in Table 3 as MT-Bench questions are open-ended. The subjectivity inherent in evaluating open-ended responses makes it more challenging to achieve consistent judgments, particularly for LLMs which may lack the nuanced understanding required to assess such responses.
6 Application
We explore the influence of LLM reliability on practical applications by focusing on how LLMs are used as judges in the Head-to-Tail benchmark in Sun et al. (2023). This benchmark aims to assess the knowledge level of LLMs by creating a set of questions about entities with varying levels of popularity ("head" - least popular, "torso" - moderately popular, "tail" - most popular). Focusing on the "head" questions, we replicate the benchmark for the Academic domain using the Vicuna-7B model.
Similar to Sun et al. (2023), we then use another LLM (Starling-LM-7B-beta, Gemma-1.1-7b-it, or Meta-Llama-3-8B-Instruct) to judge the accuracy of the benchmark’s answers, a metric termed $A_{LM}$. As we will show, this highlights reliability issues that ensue when relying on LLMs for evaluation.
6.1 Accuracy
To ensure our work is an accurate representation of Sun et al. (2023), we replicated the $A_{LM}$ results reported for the "head" dataset using the Vicuna-7B model. The $A_{LM}$ metric is calculated by the proportion of answers judged to be "correct" by the LLM judgment model. Therefore, the expected judgment responses on the accuracy of the Vicuna-7B responses are "correct" or "incorrect", with a judgment non-response considered to fall within the "incorrect" category as well. As shown in Table 3, a single instantiation of our judgment models (Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct) can closely replicate the original $A_{LM}$ values, demonstrating consistency across Sun et al. (2023) and our implementation.
$\mathbf{A_{LM}}$ Vicuna-7B
Sun et al., 2024 0.039
Starling-LM-7B-beta 0.0387
Meta-Llama-3-8B-Instruct 0.0397
Gemma-1.1-7b-it 0.0407
Table 4: The head $\mathbf{A_{LM}}$ as calculated by our judgment models Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct in a single instantiation compared to the head $\mathbf{A_{LM}}$ in Sun et al., 2023 for the Vicuna-7B model.
Figure 3 illustrates the variability in $A_{LM}$ across multiple instantiations of LLM judgments. Variation is similar across judgment models, signaling that there is no LLM judge that outperforms the other judges for this application. However, the figure reveals a notable level of variation across instantiations for all models. This variability can be attributed to inherent randomness in the LLM generation process, similar to the concept of error bars in traditional statistical analysis. In this context, the variability across instantiations results in up to a $1\%$ increase or $1\%$ decrease in the $A_{LM}$ metric, impacting the overall understanding of accuracy achieved by a model responding to the Head-to-Tail benchmark questions.
Figure 3: The variation in $A_{LM}$ of Vicuna-7B responses to Head questions across instantiations of judgment models.
6.2 Reliability
Omega Vicuna-7B
Starling-LM-7B-beta 0.9901
Meta-Llama-3-8B-Instruct 0.9896
Gemma-1.1-7b-it 0.9893
Table 5: Omega reliability of our judgment models Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct across replications for Head question responses by Vicuna-7B.
Our reliability framework sheds essential insights into the variability observed in LLM judgments. Reliability of LLM judgments across replications in this context, as seen in Table 5, is quite high. It is unsurprising that the Head-to-Tail reliability is much higher than that of our experiments given that the LLM judgments in the Head-to-Tail paradigm only have two options (’correct’ or ’incorrect’) and are provided with a ground truth in the prompt. Thus, this employment of LLM-as-a-judge is much more simplistic than our experiments and would be expected to show a higher reliability. Despite high reliability scores, Figure 3 shows that $A_{LM}$ varies across replications by ±$\pm 1\%$ as previously discussed. This underscores the potential for substantial error in downstream research that utilizes LLM-based evaluations, even when reliability is high across LLM-as-a-judge-replications. To illustrate the practical implications of this variability, consider a scenario where an LLM is used to rank different models based on their performance on a given task. A $2\%$ variation in the LLM’s judgment could lead to significant shifts in the ranking, potentially misrepresenting the true performance of the models. In the context of high-stakes applications such as ranking models to be used for medical diagnosis, even a small variation in judgment could have serious consequences. For example, an LLM-based diagnostic tool that incorrectly classifies a patient’s condition could lead to incorrect treatment and adverse health outcomes. As the stakes of LLM judgments increase, understanding and mitigating this variability becomes paramount.
7 Discussion
Current utilization of LLMs as judges often relies on single outputs as a clear delineation of an LLM’s choice in judgment, leaning upon deterministic settings (Ouyang et al., 2023; Wei et al., 2024). Using deterministic settings, however, leads to fixed randomness which masks inherent variability and fails to adequately encompass the true nature of the judgment responses, presenting a facade of reliability where one does not actually exist.
Akin to confidence intervals or standard errors typically reported in research, our proposed reliability metric for the LLM-as-a-judge paradigm provides a method for measuring uncertainty of the resulting LLM judgments and should be reported alongside LLM-as-a-judge results. To promote user understanding of model outputs, it’s crucial to maintain skepticism about the LLM’s inner workings and to be mindful of potential uncertainties, just as one would with any other model. Reliability bridges this gap. Researchers can gain a more accurate understanding of conclusions derived from LLM-as-a-judge results by reporting reliability alongside these LLM-based results.
Even for simple tasks with high reliability, LLMs can exhibit significant variability in their outputs, as illustrated in Figure 3. This variability, further emphasized by the low reliability values observed in more complex judgment tasks (Table 1), underscores the importance of acknowledging the inherent uncertainty in LLM judgments. Given the variability seen in downstream applications (Figure 3) in the presence of high reliability values (Table 5), variability across replications is particularly concerning for complex judgment setups, beyond basic correctness assessments, that are likely to be used in practice. This variability highlights the importance of informing future research with reliability. As LLMs are increasingly employed for more complex and subjective tasks, such as making critical medical decisions or assessing another LLMs’ handling of social media moderation, understanding and mitigating uncertainty becomes even more crucial. By incorporating reliability into LLM-based research, we can foster a more cautious and responsible approach to LLM usage, mitigating risks and ensuring responsible LLM employment. This includes encouraging critical evaluation of LLM judgment results and promoting transparency and accountability in LLM-based research. This will help ensure that LLMs are used effectively and ethically, maximizing their potential benefits while minimizing their risks.
8 Future Directions
Clearly, reliability is a critical factor in the successful implementation of the LLM-as-a-judge paradigm. Our work provides a foundational framework for assessing LLM evaluation reliability. To ensure responsible LLM usage, future research should develop domain-specific guidelines for interpreting reliability metrics, as different fields have varying tolerance levels for uncertainty.
Further, the impact of prompt engineering on LLM reliability has been widely discussed (Shah, 2024; Wang et al., 2024). As LLMs are susceptible to adversarial prompting (Zhu et al., 2023b), it’s crucial to assess their reliability across multiple instantiations under various prompts. This knowledge will inform improvements in prompt engineering and LLM model robustness, driving advancements in LLM development.
9 Limitations
While this work represents a significant step forward in understanding and assessing the reliability of LLMs, it is important to acknowledge its limitations. Our work is currently focused on single and multi-turn benchmarks. These benchmarks were selected to provide a robust and generalizable understanding of the impact of reliability under varying and widely used contexts, including both structured and open-ended tasks. Future research can extend our approach to additional datasets and benchmarks. Further, as discussed in Section 8, developing domain-specific practical guidelines for interpreting and applying reliability metrics is an important area for future work given that different domains have varying tolerances for variability. For instance, fields like advertising may be more tolerant of variability, while fields like medicine require much higher levels of certainty. By providing domain-specific guidance, we can ensure that LLMs are used responsibly and effectively in various applications.
Despite these limitations, our work provides a strong foundation for future research on LLM reliability. By introducing a rigorous framework for evaluating LLM-as-a-judge responses, we offer valuable insights into the impact of reliability of LLM outputs and provide a roadmap for improving the quality and trustworthiness of LLMs.
10 Ethics Statement
This research does not raise any ethical concerns based on the theories or datasets employed. The benchmarks (BBH, SQuAD, MT-Bench, and Head-to-Tail) are publicly available and utilize publicly available data sources. As such, no privacy violations are anticipated. However, researchers should exercise caution when applying this work to proprietary datasets, as these may involve specific privacy regulations and ethical considerations.